# D-69120 Heidelberg, Germany

### v2

@@ 6.4 Net torques and angular velocities
The exact difference in RAT using the canonical or the modified parametrization and, subsequently, the magnitude of the angular velocity of the grains cannot a priori be quantified due to the dependency of RATs on the spectral characteristics of the radiation field. Therefore we calculate the maximal angular velocity $\omega_{\mathrm{RAT}}$ induced by RATs using both parametrizations and compare the results to those we get using the DDSCAT data of individual grains.
@@
If the torque efficiency $Q_{\Gamma}(\lambda)$ for a single grain exhibits some sign changes over wavelength (see Fig. 4), this leads to a lower torque overall as different wavelengths exert RAT components in opposite directions. As mentioned above, the effect of positive and negative components of $Q_{\Gamma}$ canceling out appears to be small compared to the overall variation in the strength of $Q_{\Gamma}$ between different grains. The torque along $\hat{a}_{1}$ is calculated for each grain following Eq. 1, where $Q_{\Gamma}$ is used without taking the absolute value. Then, the maximal angular velocity is calculated with Eq. 3, where $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$ are also evaluated individually for each grain.
@@
Figure 8 shows the distribution of angular velocities and torques for individual SC/BA grains and the most luminous source i.e. the AGN. The distributions are broken down by grain size. We report that the resulting values of the RATs may span a range up to two orders of magnitude, with a larger spread below the mean. The mean of the distribution increases with grain size, as larger grains have higher $Q_{\Gamma}$ values in the constant part. The distributions of the different grain sizes are more distinct in $\Gamma_{\mathrm{RAT}}$ and overlap significantly for $\omega_{\mathrm{RAT}}$, due to the influence of the drag timescale (see Eq. 3).
@@
Using a parametrization describing the average value for an ensemble of grains saves the need to calculate $\omega_{\mathrm{RAT}}$ for every single grain. We therefore compare the average values obtained from individual grain calculations (shown in Fig. 8 as solid lines) with the values obtained when using the canonical parametrization for $Q_{\Gamma}$ (dotted lines). This parametrization does not model the means of the distributions well. It overestimate the values of $\Gamma_{\mathrm{RAT}}$ and $\omega_{\mathrm{RAT}}$, especially for the smaller grains.
@@
Figure 8: Histograms representing the distribution of the RAT $\Gamma_{\mathrm{RAT}}$ (left) and angular velocity $\omega_{\mathrm{RAT}}$ (right) for SC/BA grains of different grain sizes and the AGN as radiation source. Continuous lines show the mean for each distribution, dotted lines show the value calculated using the canonical parametrization (LH07). Mean values for $\omega_{\mathrm{RAT}}$ are calculated from the mean values of $\Gamma_{\mathrm{RAT}}$ shown in the lower panel as well as averages for $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$. We note that RATs by the canonical parametrization are consistently higher than the mean values of the dust aggregates.
@@
To quantify how the canonical parametrization of LH07 and the modified parametrization of our study compare with the distribution means for all combinations of radiation sources, materials, and monomer-densities, we show the distribution means of the RATs in Fig. 9 as a function the grain size $a_{\mathrm{eff}}$. The CA material in the left column, the SC material in the right column, the monomer-densities in the rows and the four radiation sources are shown with different colors in each panel. The difference between the sources is due to the linear dependence of $\Gamma_{\mathrm{RAT}}$ with energy density $u$ and anisotropy factor $\gamma$ where larger RATs are reached for stronger radiation sources (see Eq. 1 and Tab. 2). The RAT of the S1 source should be larger by a factor of 7 due to $\gamma$ and by a factor of $5.6$ due to $u$. The difference seen in Fig. 9 is larger than that with approximately 2 orders of magnitude. This is a result of the spectral energy distribution of this source. The integral in Eq. 1 puts a much larger weight on $u$ at wavelengths smaller than $\approx 1\,\mu\mathrm{m}$, as $Q_{\lambda}$ drops by about 4 orders of magnitude for larger wavelengths. This is the range where the stellar sources outshine the ISRF.
@@
In Fig. 10 we present the mean angular velocities $\omega_{\mathrm{RAT}}$ corresponding to the RATs shown in Fig. 9. Here, the variations of $\omega_{\mathrm{RAT}}$ between individual sources is different to the one of $\Gamma_{\mathrm{RAT}}$, because $\omega_{\mathrm{RAT}}$ also depends on $\tau_{\mathrm{drag}}$, which itself depends indirectly on $u$ via the temperature dependence of $\tau_{\mathrm{IR}}$. Using the dust temperatures in Tab. 3 one can estimate $\tau_{\mathrm{IR}}$ for the given temperature from Fig. 3. The gas drag timescale is independent of $u$ and about $\tau_{\mathrm{gas}}\gtrsim 10^{12}\,\mathrm{s}$ for the chosen gas density of $n_{\mathrm{gas}}=5\cdot 10^{7}\,\mathrm{m^{-3}}$. We can see, that the drag timescale is dominated by $\tau_{\mathrm{gas}}$ for the lower dust temperatures of the ISRF and S1 sources, both timescales are about equal for S2 and $\tau_{\mathrm{IR}}$ dominates for AGN, where the dust is the hottest.
Figure 9: The RATs $\Gamma_{\mathrm{RAT}}$ (solid line) per grain size $a_{\mathrm{eff}}$ calculated with the two-power-law parametrization is shown in comparison to the RATs predicted by the canonical parametrization suggested in LH07 (dotted line) as well as the average of the DDSCAT simulation data (diamonds). Each color represents one of the different sources i.e. ISRF (gray), S1 (dark blue), S2 (light blue), or AGN (red), respectively, utilizing the CA grain material (left column) and the SC material (right column) with monomer-densities BA (top row), BAM (middle row), or BAM2 (bottom row). This figure is to be compared with Fig. 8.
Figure 10: Same as Fig. 9 but for angular velocity $\omega_{\mathrm{RAT}}$. We emphasize that the canonical parametrization may over-predict the magnitude of $\omega_{\mathrm{RAT}}$ up to one order of magnitude higher for BAM grain aggregates as the two-power-law parametrization.
@@
Therefore, the difference in $\omega_{\mathrm{RAT}}$ depends only on $\Gamma_{\mathrm{RAT}}$ for the ISRF and S1, but includes $\tau_{\mathrm{IR}}$ for S2 and AGN. This is why AGN is much closer to S2 in $\omega_{\mathrm{RAT}}$ ($1\,\mathrm{dex}$) than in $\Gamma_{\mathrm{RAT}}$ ($3\,\mathrm{dex}$). Even though the torque is much larger for the AGN source, the grains do not rotate that much faster, because they also get much hotter, increasing the IR-drag. The difference in $\tau_{\mathrm{IR}}$ between the two materials also leads to smaller angular velocity for S2 and AGN for SC compared to CA, but no difference between the two materials for S1 and ISRF, as $\tau_{\mathrm{gas}}$ dominates there. The difference between monomer-densities is a decrease in $\Gamma_{\mathrm{RAT}}$ from BA to BAM2 smaller than one order of magnitude, due to $Q_{\Gamma}$ decreasing with monomer-density. The resulting decrease with monomer-density in $\omega_{\mathrm{RAT}}$ is smaller, due to $I_{\mathrm{max}}$ also decreasing with monomer-density.
@@
As expected RATs have a strong scaling with $a_{\mathrm{eff}}$ and it does not change much between sources, materials and monomer-densities. However, for the angular velocity $\omega_{\mathrm{RAT}}$, the results are less uniform as shown in Fig. 10. Here, there is no scaling with grain size for the ISRF source, and a weak one for the AGN, which flattens towards the larger grains. For the stellar sources $\omega_{\mathrm{RAT}}$ first increases and then flattens for SC, even decreasing towards larger grains for S1. In the case of CA $\omega_{\mathrm{RAT}}$ even scales inversely with grain size for the stellar sources. We remark that the scaling of $\omega_{\mathrm{RAT}}$ with grain size depends on the interplay between multiple factors: spectral shape of the source, scaling of the drag timescale, which depends on the material and structure of the grain, and whether $\tau_{\mathrm{IR}}$ or $\tau_{\mathrm{gas}}$ dominates for the given parameters. As outlined in the pioneering work of Hoang et al. (2019) and Hoang (2019) RATs may spin-up dust grains up to a characteristic angular velocity $\omega_{\mathrm{disr}}$ where the internal grain structure can no longer resist the disruption because of centrifugal forces (RATD). The predicted order of magnitude of for RATD to occur is $\omega_{\mathrm{disr}}\approx 10^{9}\mathrm{s}^{-1}$. Later, ab-initio numerical N-body simulations of the disruption of rotating BA aggregates revealed that aggregates larger than $a_{\mathrm{eff}}=300\ \mathrm{nm}$ are more resistant against RATD because of deformation. We emphasize that the most luminous sources, AGN and S2, respectively, may spin up the dust aggregates beyond the limit of $\omega_{\mathrm{RAT}}>\omega_{\mathrm{disr}}$ (see Fig. 10) even though RATs are less efficient for ballistic aggregates. Hence, most of the grains may become disrupted by RATD in proximity of an AGN or S2-like star. So ballistic aggregates are more likely to grow deep inside of dense molecular clouds where they are shielded from a strong radiation field (Bethell et al. 2007).
@@
As presented in Fig. 10, the canonical parametrization gives more accurate results for the larger grain aggregates, as they experience the strongest torques. We point out that overall the canonical parametrization overestimates the mean angular velocity $\omega_{\mathrm{RAT}}$, with differences up to one order of magnitude. In comparison, the two-power-law parametrization resembles the mean of the DDSCAT data more precisely and performs equally well for all sources, materials, monomer-densities, and sizes.
@@
A way to quantify the efficiency of RATs of porous aggregates was proposed by Tatsuuma & Kataoka (2021), TK21 hereafter. The TK21 parametrization takes the porosity (monomer-density) of the grain aggregates into account via the volume filling factor $\phi$, where $\phi=1$ represents a solid grain and $\phi=0$ represents completely empty space. Here, the volume filling factor $\phi$ is added into the canonical parametrization as follows
@@
Fig. 11 shows the average angular velocity $\omega_{\mathrm{RAT}}$ in comparison with the TK21 parametrization and canonical parametrization of LH07. The model proposed by TK21 clearly improves upon the canonical parametrization of LH07 for the less porous BAM1 and BAM2 grains, but now slightly under-predicts $\omega_{\mathrm{RAT}}$ for the weaker sources in BAM1. For the more porous BA grains however TK21 clearly under-predicts $\omega_{\mathrm{RAT}}$, as this is the regime, where the canonical parametrization performed best. As shown in Fig. 11, merely introducing $\phi$ to take the grain porosity into account does not bring the canonical parametrization into agreement with our results, indicating that complex grain shapes differ in their interaction with radiation compared to compact grains in more fundamental ways than just by making the grains porous.
Consequently we may summarize, that even though there is a large spread in $\omega_{\mathrm{RAT}}$ between individual grains (see Fig. 8) due to random variation in grain shape, using a parametrization of $Q_{\Gamma}$ is still valid to model the average $\omega_{\mathrm{RAT}}$ for an ensemble of ballistic grain aggregates.
For the three parametrizations shown in this work the two-power-law parametrization is the most accurate choice for this task.
Figure 11: The same as Fig. 10 but showing the canonical parametrization of LH07 (dotted lines) in comparison with the parametrization for porous grains suggested in TK21 (dashed lines).
@@ 7 Discussion and Caveats
An underlying assumption in our RAT calculations using grain aggregates is the anisotropy of incoming radiation. As DDSCAT only simulates anisotropic radiation, we used the anisotropy factor $\gamma$ to treat a certain percentage of the source radiation as anisotropic. For instance a grain in a molecular cloud around a forming star receives a portion $\gamma$ of the stellar radiation directly (anisotropically) and the rest isotropically due to scattering on surrounding particles. Such processes are not fully accounted for in this work, however, we do not expect torques from isotropic radiation to dominate torques from anisotropic radiation, especially for our grain model with no variations of optical properties across the grains surface.
@@
We also assume perfect alignment of the grain rotation axis with the direction of the radiation i.e. $\Theta=0$. This is based on the fact, that grains preferentially align with the direction of incoming radiation in the absence of magnetic fields (see LH07). Though magnetic fields are present in the typical environments probed here and the grains may not reach suprathermal angular velocities. For a more detailed picture, one would need DDSCAT simulations sampling the whole range of $\Theta$-angles and, subsequently, be used to calculate phase space trajectories of grains in $\Theta$ and $\omega$, as well as including typical magnetic fields. Such an approach may allow to verify if the assumption of fixed orientation was justified, and if not, to gauge how large the deviation is. However, this is beyond the scope of our current study.
@@
Using a parametrization to model the average angular velocity is still possible with the more detailed grain model, even including positive and negative values of $Q_{\Gamma}(\lambda)$ for individual grains. If needed it might also be possible to parametrize the distribution of $\omega_{\mathrm{RAT}}$ values (see Fig. 8), or other quantities using larger grain ensembles.
@@
As we were limited to a small range of grain sizes due to the computational restrictions of DDSCAT, future simulation setups should be designed to test if the two-power-law parametrization for $|Q_{\Gamma}|$ also holds for larger grains. If this is the case and the two-power-law shape is also seen for general orientations of the grain, it would stand to reason, that the form of $|Q_{\Gamma}|$ following two power-laws is intrinsic to the nature of grains composed of monomers, like the single power-law of the canonical parametrization is for compact grains.
It might be convenient to generalize the parameters of the parametrization and give them a functional form dependent on monomer-density and grain size. Fig. 12 shows the constant part $c$ of the parametrization as a function of grain size for different monomer-densities. We can see that the linear increase of $c$ with grain size has a different slope for each monomer-density, with the slope roughly doubling from BAM2 to BAM1 and tripling from BAM2 to BA. It is therefore possible, that for each material there is a functional form of $c$ as a linear function with the slope depending on monomer-density. To see if and how the exponents $q_{1}$ and $q_{2}$ depend on material or monomer-density would require a larger sample of monomer-densities and refractive indices (materials).
We have shown, that aggregated grains have a different functional form of $Q_{\Gamma}(\lambda)$ than compact grains and therefore do not reach as high angular velocities as compact grains would in the same circumstances, with differences of up to an order of magnitude for the smallest grains. However we see that our results and the canonical parametrization agree for the least dense and largest grains (BA, $250\,\mathrm{nm}$). If the trend of increasing $c$ for larger grain size and lower monomer-density (see Fig. 12) holds true, we would expect $\Gamma_{\mathrm{RAT}}$ to further increase for less dense and larger grains. The relation of lower monomer-density leading to higher torques makes intuitive sense considering it increases the total surface area of the grain and therefore the cross section for interaction with radiation just as increasing the grain size does. While the torque increases, $\omega_{\mathrm{RAT}}$ however does not increase further for larger grains and in some cases even decreases again for the largest grains. Aggregate grains larger than the ones simulated here may potentially not reach much higher angular velocities, as the drag also increases with size.
How much $\omega_{\mathrm{RAT}}$ increases for higher radiation energy densities strongly depends on the optical properties and therefore on the composition of the grains, due to different drag-effects dominating in different environments. The significantly increased surface area and more complex geometry of aggregated grains has to be taken into account.
In environments with highly ionized gas and charged grains, such as an AGN, plasma drag in addition to gas drag and photon drag might additionally play a role for grain dynamics (Draine & Lazarian 1998). However, the charging of BA dust adds an additional level of complexity where individual monomers may carry different charges and even undergo some stochastic fluctuation over time (see e.g. Matthews et al. 2013; Ma et al. 2013; Matthews et al. 2018). Including the plasma drag in the context of BA goes beyond the scope of this paper. However, the difference of the total grain charge of a BA compared to an equivalent sphere should be less than one order of magnitude as shown in Ma et al. (2013) and Matthews et al. (2018). Hence, we estimate that the canonical parametrization of RAT would still overestimate the resulting angular velocity $\omega_{\mathrm{RAT}}$ for ballistic aggregates in proximity of an AGN. The dust grains were assumed to have no relative velocity to the gas, but it is possible, that grains close to strong radiation fields experiencing linear acceleration by the radiation pressure will reach high drift velocities relative to the surrounding gas. We emphasize that the resulting rotation of the BA grain would then also be influenced by mechanical torques (METs, see e.g. Lazarian & Hoang 2007b; Das & Weingartner 2016; Hoang et al. 2018; Reissl et al. 2023) working in a similar way to RATs, with gas collisions instead of photon collisions. However, the question of how METs may enhance the grain rotation in the presence of RATs is still a matter of debate (see also Lazarian & Hoang 2021).
We emphasize once again that our modeling of RATs does not include the dynamical processes associated with the presence of an external magnetic field. As noted in LH07 for an analytical RAT model, the appearance of attractors in the $\Theta-\omega$ phase space is tightly connected to the ratio ${q=\max(Q_{\mathrm{e1}})/\max(Q_{\mathrm{e2}})}$ when an external field is present. Later, this finding was confirmed by Herranen et al. (2019) and Herranen (2020) based on grain shapes modeled as Gaussian ellipsoids. We note, however, that a similar parameter $q$ for pure mechanical alignment was not found (Das & Weingartner 2016; Reissl et al. 2023). Hence, it remains to be seen if the ratio $q$ is a reliable proxy to predict the long-term stable alignment with the direction of the magnetic field, considering RATs as a spin-up mechanism of ballistic aggregates. Moreover, in the presence of a magnetic field, this is not only a matter of the efficiency of the RATs. Due to internal dissipation processes in paramagnetic and ferromagnetic materials, a fraction of the angular velocity $\omega$ is transferred into internal heat (see e.g. Davis & Greenstein 1951; Dolginov & Mytrophanov 1976; Purcell 1979). Consequently, the $\omega$ reported in our study may be lower, depending on the magnetic field strength and grain orientation. A similar effect of a reduction in $\omega$ for METs is shown by Reissl et al. (2023). However, to what degree a magnetic field would reduce the maximal possible angular velocity $\omega_{\mathrm{RAT}}$ of BAM aggregates remains to be addressed in forthcoming publications.

### v3

@@ 6.4 Net torques and angular velocities
The exact difference in RAT using the canonical or the modified parameterization, and subsequently, the magnitude of the angular velocity of the grains, cannot a priori be quantified because the RATs depend on the spectral characteristics of the radiation field. Therefore, we calculated the maximum angular velocity $\omega_{\mathrm{RAT}}$ induced by RATs using both parameterizations and compared the results to those we obtained using the DDSCAT data of individual grains.
@@
If the torque efficiency $Q_{\Gamma}(\lambda)$ for a single grain exhibits some sign changes over wavelength (see Fig. 4), this leads to a lower torque overall as different wavelengths exert RAT components in opposite directions. As mentioned above, the effect of positive and negative components of $Q_{\Gamma}$ canceling out appears to be small compared to the overall variation in the strength of $Q_{\Gamma}$ between different grains. The torque along $\hat{a}_{1}$ was calculated for each grain following Eq. 1, where $Q_{\Gamma}$ was used without taking the absolute value. Then, the maximum angular velocity was calculated with Eq. 3, where $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$ were also evaluated individually for each grain.
@@
Figure 8 shows the distribution of angular velocities and torques for individual SC BA grains and the most luminous source, that is, the AGN. The distributions are broken down by grain size. We report that the resulting values of the RATs may span a range of up to two orders of magnitude, with a larger spread below the mean. The mean of the distribution increases with grain size because larger grains have higher $Q_{\Gamma}$ values in the constant part. The distributions of the different grain sizes are more distinct in $\Gamma_{\mathrm{RAT}}$ and overlap significantly for $\omega_{\mathrm{RAT}}$ because of the drag timescale (see Eq. 3).
@@
A parameterization that describes the average value for an ensemble of grains means that we do not need to calculate $\omega_{\mathrm{RAT}}$ for every single grain. We therefore compared the average values obtained from individual grain calculations (shown in Fig. 8 as solid lines) with the values obtained for the canonical parameterization for $Q_{\Gamma}$ (dotted lines). This parameterization does not model the means of the distributions well. It overestimate the values of $\Gamma_{\mathrm{RAT}}$ and $\omega_{\mathrm{RAT}}$, especially for the smaller grains.
@@
Figure 8: Histograms representing the distribution of the RAT $\Gamma_{\mathrm{RAT}}$ (left) and angular velocity $\omega_{\mathrm{RAT}}$ (right) for SC BA grains of different sizes and the AGN as the radiation source. The continuous lines show the mean for each distribution, and the dotted lines show the value calculated using the canonical parameterization (LH07). The mean values for $\omega_{\mathrm{RAT}}$ are calculated from the mean values of $\Gamma_{\mathrm{RAT}}$ shown in the lower panel as well as averages for $\tau_{\mathrm{drag}}$ and $I_{\mathrm{max}}$. We note that the RATs by the canonical parameterization are consistently higher than the mean values of the dust aggregates.
@@
To quantify how the canonical parameterization of LH07 and the modified parameterization of our study compare with the distribution means for all combinations of radiation sources, materials, and monomer densities, we show the distribution means of the RATs in Fig. A as a function the grain size $a_{\mathrm{eff}}$. The CA material in the left column, the SC material in the right column, the monomer densities in the rows, and the four radiation sources are shown with different colors in each panel. The difference between the sources is due to the linear dependence of $\Gamma_{\mathrm{RAT}}$ on the energy density $u$ and anisotropy factor $\gamma$, where larger RATs are reached for stronger radiation sources (see Eq. 1 and Tab. 2). The RAT of the S1 source should be larger by a factor of 7 due to $\gamma$ and by a factor of $5.6$ due to $u$. The difference seen in Fig. A is larger than that with approximately two orders of magnitude. This is a result of the spectral energy distribution of this source. The integral in Eq. 1 places a much larger weight on $u$ at wavelengths shorter than $\approx 1\,\mu\mathrm{m}$, as $Q_{\lambda}$ drops by about four orders of magnitude for longer wavelengths. In this range, the stellar sources outshine the ISRF.
@@
In Fig. 10 we present the mean angular velocities $\omega_{\mathrm{RAT}}$ corresponding to the RATs shown in Fig. A. Here, the variations in $\omega_{\mathrm{RAT}}$ between individual sources are different from the variation in $\Gamma_{\mathrm{RAT}}$ because $\omega_{\mathrm{RAT}}$ also depends on $\tau_{\mathrm{drag}}$, which itself depends indirectly on $u$ via the temperature dependence of $\tau_{\mathrm{IR}}$. Using the dust temperatures in Tab. 3, we can estimate $\tau_{\mathrm{IR}}$ for the given temperature from Fig. 3. The gas-drag timescale is independent of $u$ and about $\tau_{\mathrm{gas}}\gtrsim 10^{12}\,\mathrm{s}$ for the chosen gas density of $n_{\mathrm{gas}}=5\cdot 10^{7}\,\mathrm{m^{-3}}$. The drag timescale is dominated by $\tau_{\mathrm{gas}}$ for the lower dust temperatures of the ISRF and S1 sources, both timescales are about equal for S2, and $\tau_{\mathrm{IR}}$ dominates for the AGN, where the dust is hottest.
@@
Therefore, the difference in $\omega_{\mathrm{RAT}}$ only depends on $\Gamma_{\mathrm{RAT}}$ for the ISRF and S1, but it includes $\tau_{\mathrm{IR}}$ for S2 and AGN. The AGN therefore is much closer to S2 in $\omega_{\mathrm{RAT}}$ ($1\,\mathrm{dex}$) than in $\Gamma_{\mathrm{RAT}}$ ($3\,\mathrm{dex}$). Even though the torque is much stronger for the AGN source, the grains do not rotate that much faster because they also become much hotter, which increases the IR drag. The difference in $\tau_{\mathrm{IR}}$ between the two materials also leads to a lower angular velocity for S2 and AGN for SC compared to CA, but there is no difference between the two materials for S1 and ISRF because $\tau_{\mathrm{gas}}$ dominates there. The difference between monomer densities is a decrease in $\Gamma_{\mathrm{RAT}}$ from BA to BAM2 that is smaller than one order of magnitude because $Q_{\Gamma}$ decreases with monomer density. The resulting decrease with monomer density in $\omega_{\mathrm{RAT}}$ is smaller because $I_{\mathrm{max}}$ also decreases with monomer density.
@@
As expected, RATs strongly scale with $a_{\mathrm{eff}}$, and this does not change much between sources, materials, and monomer densities. However, for the angular velocity $\omega_{\mathrm{RAT}}$, the results are less uniform, as shown in Fig. 10. Here, there is no scaling with grain size for the ISRF source and a weak scaling for the AGN, which flattens toward the larger grains. For the stellar sources, $\omega_{\mathrm{RAT}}$ first increases and then flattens for SC, and it even decreases toward larger grains for S1. In the case of CA, $\omega_{\mathrm{RAT}}$ even scales inversely with grain size for the stellar sources. We remark that the scaling of $\omega_{\mathrm{RAT}}$ with grain size depends on the interplay between multiple factors: The spectral shape of the source and the scaling of the drag timescale, which depends on the material and structure of the grain, and whether $\tau_{\mathrm{IR}}$ or $\tau_{\mathrm{gas}}$ dominates for the given parameters. As outlined in the pioneering work of Hoang et al. (2019) and Hoang (2019), RATs may spin up dust grains to a characteristic angular velocity $\omega_{\mathrm{disr}}$, where the internal grain structure can no longer resist the disruption because of centrifugal forces (RATD). The predicted order of magnitude for RATD to occur is $\omega_{\mathrm{disr}}\approx 10^{9}\mathrm{s}^{-1}$. Later, ab initio numerical N-body simulations of the disruption of rotating BA aggregates revealed that aggregates larger than $a_{\mathrm{eff}}=300\ \mathrm{nm}$ are more resistant against RATD because they are deformed. We emphasize that the most luminous sources, AGN and S2 may spin up the dust aggregates beyond the limit of $\omega_{\mathrm{RAT}}>\omega_{\mathrm{disr}}$ (see Fig. 10), even though RATs are less efficient for ballistic aggregates. Hence, most of the grains may become disrupted by RATD in the proximity of an AGN or S2-like star. This means that ballistic aggregates are more likely to grow deep inside of dense molecular clouds, where they are shielded from a strong radiation field (Bethell et al. 2007).
@@
As presented in Fig. 10, the canonical parameterization gives more accurate results for the larger grain aggregates because they experience the strongest torques. We point out that the canonical parameterization overestimates the mean angular velocity $\omega_{\mathrm{RAT}}$ overall, with differences of up to one order of magnitude. In comparison, the two-power-law parameterization resembles the mean of the DDSCAT data more precisely and performs equally well for all sources, materials, monomer densities, and sizes.
@@
A way to quantify the efficiency of RATs of porous aggregates was proposed by Tatsuuma & Kataoka (2021) (TK21 hereafter). The TK21 parameterization takes the porosity (monomer density) of the grain aggregates into account via the volume filling factor $\phi$, where $\phi=1$ represents a solid grain, and $\phi=0$ represents completely empty space. Here, the volume-filling factor $\phi$ was added to the canonical parameterization as follows:
@@
Fig. 11 shows the average angular velocity $\omega_{\mathrm{RAT}}$ in comparison with the TK21 parameterization and canonical parameterization of LH07. The model proposed by TK21 clearly improves upon the canonical parameterization of LH07 for the less porous BAM1 and BAM2 grains, but now slightly underpredicts $\omega_{\mathrm{RAT}}$ for the weaker sources in BAM1. For the more porous BA grains, however, TK21 clearly underpredicts $\omega_{\mathrm{RAT}}$ because in this regime, the canonical parameterization performed best. As shown in Fig. 11, merely introducing $\phi$ to take the grain porosity into account does not bring the canonical parameterization into agreement with our results. This indicates that complex grain shapes differ in their interaction with radiation compared to compact grains in more fundamental ways than just by making the grains porous.
Consequently, we may summarize that even though there is a large spread in $\omega_{\mathrm{RAT}}$ between individual grains (see Fig. 8) due to random variation in the grain shape, it is still valid to use a parameterization of $Q_{\Gamma}$ to model the average $\omega_{\mathrm{RAT}}$ for an ensemble of ballistic grain aggregates.
For the three parameterizations shown in this work, the two-power-law parameterization is the most accurate choice for this task.
@@ 7 Discussion and caveats
An underlying assumption in our RAT calculations using grain aggregates is the anisotropy of incoming radiation. As DDSCAT only simulates anisotropic radiation, we used the anisotropy factor $\gamma$ to treat a certain percentage of the source radiation as anisotropic. For instance, a grain in a molecular cloud around a forming star receives a portion $\gamma$ of the stellar radiation directly (anisotropically) and the rest isotropically from scattering on surrounding particles. These processes are not fully accounted for in this work, but we do not expect torques from isotropic radiation to dominate torques from anisotropic radiation, especially for our grain model without variations in the optical properties across the grain’s surface.
@@
We also assumed perfect alignment of the grain rotation axis with the direction of the radiation, that is, $\Theta=0$. This is based on the fact that grains preferentially align with the direction of incoming radiation in the absence of magnetic fields (see LH07), even though magnetic fields are present in the typical environments probed here and the grains may not reach suprathermal angular velocities. For a more detailed picture, DDSCAT simulations that sample the whole range of $\Theta$ angles are required that are subsequently used to calculate phase-space trajectories of grains in $\Theta$ and $\omega$, and they would also need to include typical magnetic fields. This approach may allow us to verify whether the assumption of a fixed orientation was justified, and if not, to gauge how large the deviation is. However, this is beyond the scope of our current study.
@@
It is still possible to use a parameterization to model the average angular velocity with the more detailed grain model, even including positive and negative values of $Q_{\Gamma}(\lambda)$ for individual grains. If needed, it might also be possible to parameterize the distribution of $\omega_{\mathrm{RAT}}$ values (see Fig. 8), or other quantities using larger grain ensembles.
@@
As we were limited to a small range of grain sizes due to the computational restrictions of DDSCAT, future simulation setups should be designed to test whether the two-power-law parameterization for $|Q_{\Gamma}|$ also holds for larger grains. If this is the case and the two-power-law shape is also seen for general orientations of the grain, it would stand to reason that the form of $|Q_{\Gamma}|$ following two power laws is intrinsic to the nature of grains composed of monomers, in the same way as the single power law of the canonical parameterization is for compact grains.